Among 157 African-American men, the mean systolic blood pressure was 146 mm Hg with a standard deviation of 27. We wish to know if on the basis of these data, we may conclude that the mean systolic blood pressure for a population of African-American is 140.



1
Expert's answer
2021-04-15T06:48:56-0400

n=157 , sample mean=146, s=27, α=0.01

"Z = \dfrac{{\overline{X}-\mu_0}}{\frac{s}{\sqrt{n}}}=\dfrac{146-140}{\frac{27}{\sqrt{157}}} = 2.78"

We reject H0 if Z>Z1-α = Z0.99= 2.33

Hence, we can conclude that the mean systolic blood pressure for a population of African-American is greater than 140 mm Hg.